OpenStudy (anonymous):

Segments GJ and JC are the same length. Find GJ and JC to the nearest tenth of a centimeter. (Enter only the number.)

OpenStudy (anonymous):

Segments GJ and JC are the same length. Find GJ and JC to the nearest tenth of a centimeter. (Enter only the number.)

OpenStudy (anonymous):

KC = 35 mm right?

OpenStudy (hoblos):

36.9 cm

OpenStudy (anonymous):

Yes jimmyrep!

OpenStudy (anonymous):

36.9 is correct by pythagors AC = sqrt(50^2 + 150^2) = 158.11 JC = (158.11 * 35 )/ 150 = 36.9